A building is casting a 100 ft. shadow at the same time that a 5 ft. post is casting a 1 ft. shadow.
Write a proportion that involves the height of the building, the height of the post, and legnths of the shadows.
How high is the building?

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

let x="height of building" ___ x/100=5/1

Answer by stanbon(75887)

A building is casting a 100 ft. shadow at the same time that a 5 ft. post is casting a 1 ft. shadow.
Write a proportion that involves the height of the building, the height of the post, and length of the shadows.
---------
Let height of the building be "x":
Proportion:
x/100 = 5/1
x = 500 ft
------------------------
How high is the building?
The building is 500 ft. high
======================
Cheers,
Stan H.